Train A and Train B depart at the same time from
the same station for a destination 4,000 miles away.
If Train A travels at constant rate of 170 miles per
hour and Train B travels at a constant rate of 200
miles per hour, how far will the two trains be from
each other, in miles, when Train B reaches its destination?

Answer by ewatrrr(24785)


Hi
D = r*t or t = D/r
Train B reaches its destination in:
4000mi/200mph = 20h
Distance Train A will have traveled in 20hr
170*20 = 3400mi
At this time they will be (4000mi-3400mi) or 600mi apart